# Certified Defense on the Fairness of Graph Neural Networks

### v2

@@ B.6 Ordering the Inner and Outer Defense
Figure 5: An example illustrating how ELEGANT works with a different order to achieve certified defense.
@@
We first review the general pipeline to achieve certified fairness defense. Specifically, we first model the fairness attack and defense by formulating the bias indicator function $g$. Then, we achieve certified defense over the node attributes for $g$, which leads to classifier $\tilde{g}_{\bm{X}}$. Finally, we realize certified defense for $\tilde{g}_{\bm{X}}$ over the graph topology, which leads to classifier $\tilde{g}_{\bm{A},\bm{X}}$. In general, we may consider the certified defense over node attributes and graph topology as the inner certified classifier and outer certified classifier, respectively. Now, a natural question is: is it possible to achieve certified defense in a different order, i.e., first achieve certified defense over the graph topology (as the inner classifier), and then realize certified defense over the node attributes (as the outer classifier)? Note that this is not the research focus of this paper, but we will provide insights about this question. In fact, it is also feasible to achieve certified defense in the reversed order compared with the approach presented in our paper. We provide an illustration in Figure 5. We follow a similar setting to plot this figure as in Section 3.3. Specifically, in case (1), both $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$ lead to a positive outcome for $g$; in case (2), both $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$ lead to a negative outcome. However, considering the Gaussian distribution around $\bm{X}_{i,j}$, samples will fall around case (1) with a much higher number compared with case (2). Hence, in this example, it would be reasonable to assume that the classifier with Bernoulli noise over graph topology (the inner certified classifier) will return 1 with a higher probability. This example thus illustrates how certification following a different order returns 1.
@@
However, such a formulation bears higher computational costs in calculating the certified budgets. The reason is that we are able to utilize a closed-form solution to calculate $\epsilon_{\bm{X}}$ based on a set of Gaussian noise and the corresponding output from the bias indicator function. However, based on a set of Bernoulli noise and the corresponding output from the bias indicator function, we will need to solve the optimization problem given in Theorem 2 to calculate $\epsilon_{\bm{A}}$, which bears a higher time complexity than calculating $\epsilon_{\bm{X}}$. If we follow the strategy provided in Section 3.4 to calculate the inner and outer certification budgets, the certified budget of the inner certification will always be calculated multiple times, while the certified budget of the outer certification will only be calculated once. Considering the high computational cost of calculating $\epsilon_{\bm{A}}$, we thus argue that it is more efficient to realize the certification over graph topology as the outer certified classifier.
B.7 Time Complexity Analysis
@@
Experimental. We perform a study of running time, and we present the results in Table 4. Specifically, we compare the running time of a successful certification under 30,000 random noise samples and a regular training-inference cycle with vanilla GCN. We observe that (1) although ELEGANT improves the computational cost compared with the vanilla GNN backbones, the running time remains acceptable; and (2) ELEGANT has less running time growth rate on larger datasets. For example, E-SAGE has around 10x running time on German Credit (a smaller dataset) while only around 4x on Credit Default (a larger dataset) compared to vanilla SAGE. Hence we argue that ELEGANT bears a high level of usability in terms of complexity and running time.
B.8 Additional Results on Different GNN Backbones & Baselines
@@
We perform additional experiments over two popular GNNs, including APPNP [33] and GCNII [7], to evaluate the generalization ability of ELEGANT onto different backbones. We present all numerical results in Table 5 (in terms of accuracy), Table 6 (in terms of fairness), and Table 7 (in terms of FCR). We observe that ELEGANT achieves comparable utility, a superior level of fairness, and a large percentage of FCR. This verifies the satisfying usability of ELEGANT, which remains consistent with the paper.
@@
In addition, we provide a detailed fairness comparison between ELEGANT and robust GNNs from [24] and [48] in Table 8. We observe that the best performances still come from the GNNs equipped with ELEGANT on all datasets. Hence we argue that ELEGANT exhibits satisfying performance in usability, which remains consistent with the discussion in the paper.
@@
Why ELEGANT Improves Fairness? We note that improving fairness is a byproduct of ELEGANT, and our focus is to achieve certification over the fairness level of the prediction results. We now provide a detailed discussion about why fairness is improved here. First, existing works found that the distribution difference in the node attribute values and edge existence across different subgroups is a significant source of bias [11, 9, 15]. However, adding noise on both node attributes and graph topology may reduce such distributional divergence and mitigate bias. Second, As mentioned in Section 3.4, the proposed strategy to obtain the output predictions in ELEGANT is to select the fairest result among the output set $\hat{\mathcal{Y}}^{\prime}$, where each output is derived based on a sample $\bm{\Gamma}_{\bm{A}}^{\prime}\in\bar{\mathcal{A}}^{\prime}$ (i.e., $\mathrm{argmin}_{\hat{\bm{Y}}^{\prime}}\pi(\hat{\bm{Y}}^{\prime},\mathcal{V}_{\text{tst}})\;\;s.t.\;\hat{\bm{Y}}^{\prime}\in\hat{\mathcal{Y}}^{\prime}$). Such a strategy provides a large enough probability to achieve certification in light of Proposition 1. Meanwhile, we point out that such a strategy also helps to significantly improve fairness since highly biased outputs are excluded.
B.9 Complementary Results
@@
We provide the results in terms of $\Delta_{EO}$ for Table 1 in Table 9, and we present the results of the baselines for Figure 2 in Table 10, Table 11, Table 12, and Table 13. For Table 9, we observe that ELEGANT does not constantly show a lower value of $\Delta_{EO}$. This is because the certification goal in Table 1 is $\Delta_{SP}$ instead of $\Delta_{EO}$. In addition, we note that debiasing existing GNN models is not the goal of this paper. In addition, we provide the corresponding results in terms of accuracy for Figure 3 in Table 14 and Table 15. We observe that although most performance remains stable, a stronger noise (i.e., larger $\sigma$ and smaller $\beta$) generally leads to worse but still comparable performance. This is consistent with the discussion in Section 4.4, and this has been taken into consideration in the discussion of the parameter selection strategy in Section 4.4.
@@
Table 4: Comparison of running time (in seconds) on different datasets using different methods.
@@
Table 5: Performance comparison of classification accuracy. Numbers are in percentage.
@@
Table 6: Comparison of fairness (measured with $\Delta_{SP}$). Numbers are in percentage.
@@
Table 7: Performance in FCR on different datasets and backbone GNNs. Numbers are in percentage.
@@
Table 8: Comparison of fairness (measured with $\Delta_{SP}$). Numbers are in percentage.
@@
       [24]        14.8 $\pm$ 18.3        9.59 $\pm$ 0.65        3.84 $\pm$ 0.17
       [48]        3.66 $\pm$ 0.52        8.04 $\pm$ 2.97        7.10 $\pm$ 5.10
Table 9: The $\Delta_{EO}$ of Table 1 in the paper. All numerical numbers are in percentage.
@@
Table 10: The results under ($2^{0}$, $10^{-1}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ Figure 2. All numerical numbers are in percentage.
@@
Table 11: The results under ($2^{1}$, $10^{0}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ Figure 2. All numerical numbers are in percentage.
@@
Table 12: The results under ($2^{2}$, $10^{1}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ for Figure 2. All numerical numbers are in percentage.
@@
Table 13: The results under ($2^{3}$, $10^{2}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ for Figure 2. All numerical numbers are in percentage.
@@
Table 14: Classification accuracy in Figure 3(a) with different settings. Numbers are in percentage.
@@
Table 15: Classification accuracy in Figure 3(b) with different settings. Numbers are in percentage.
@@ Appendix C Additional Discussion
C.1 Why Certify A Classifier on top of An Optimized GNN?
@@
C.2 What Is the Difference Between the Attacking Performance of GNNs and the Fairness of GNNs?
C.3 Certification Without Considering the Binary Sensitive Attribute
C.4 How Do the Main Theoretical Findings Differ From Existing Works on Robustness Certification of GNNs on Regular Attacks?
Most existing works for robustness certification can only defend against attacks on either node attributes or graph structure. Due to the multi-modal input data of GNNs, existing works usually fail to handle the attacks over node attributes and graph structure at the same time. However, ELEGANT is able to defend against attacks over both data modalities. This necessitates using both continuous and discrete noises for smoothing and the analysis for joint certification in the span of the two input data modalities (as shown in Figure 1).
C.5 Discussion: Difference with Existing Similar Works
Here we mainly focus on discussing the difference between this work and [4].
We note that (1) the randomized smoothing technique adopted in [4] is different from the proposed randomized smoothing approach on the graph topology in this paper and (2) the techniques in [4] tackle a different problem from this paper. We elaborate on more details below.
The techniques in [4] are different from this paper. Although both randomized smoothing approaches are able to handle binary data, we note that the randomized smoothing approach proposed in [4] is data-dependent. However, the proposed randomized smoothing approach in this paper is data-independent. We note that in practice, a data-independent approach enables practitioners to pre-generate noises, which significantly improves usability.
The studied problem in [4] is different from this paper. Although the authors claimed to achieve a joint certificate for graph topology and node attributes in [4], all node attributes are assumed to be binary, which can only be applied to cases where these attributes are constructed as bag-of-words representations (as mentioned in the second last paragraph in the Introduction of [4]). However, in this work, we follow a more realistic setting where only graph topology is assumed to be binary while node attributes are considered as continuous. This makes the problem more difficult to handle, since different strategies should be adopted for different data modalities. In summary, compared with [4], the problem studied in this paper is more realistic and more suitable for GNNs.
Based on the discussion above, we would like to note that no existing work can be directly adopted to tackle the studied problem in this paper.
C.6 Additional Experiments on Different Datasets
To further validate the performance of the proposed method, we also perform experiments with the same commonly used popular GNN backbone models (as Section 4.2) on two Pokec datasets, namely Pokec-z and Pokec-n. We present the experimental results in Table 16, where all numerical numbers are in percentage. We observe that (1) the GNNs equipped with ELEGANT achieve comparable node classification accuracy; (2) the GNNs equipped with ELEGANT achieve consistently lower levels of bias; and (3) the values of the Fairness Certification Rate (FCR) for all GNNs equipped with ELEGANT exceed 90%, exhibiting satisfying usability. All three observations are consistent with the experimental results and the corresponding discussion presented in Section 4.2. Therefore, we argue that the effectiveness of the proposed approach is not determined by the dataset and is well generalizable over different graph datasets.
Table 16: Experimental results on Pokec-z and Pokec-n datasets.
@@ C.6 Additional Experiments on Different Datasets
C.7 Scalability of ELEGANT

### v3

@@ B.4. Ordering the Inner and Outer Defense
Figure 5. An example illustrating how ELEGANT works with a different order to achieve certified defense.
@@
We first review the general pipeline to achieve certified fairness defense. Specifically, we first model the fairness attack and defense by formulating the bias indicator function $g$. Then, we achieve certified defense over the node attributes for $g$, which leads to classifier $\tilde{g}_{\bm{X}}$. Finally, we realize certified defense for $\tilde{g}_{\bm{X}}$ over the graph topology, which leads to classifier $\tilde{g}_{\bm{A},\bm{X}}$. In general, we may consider the certified defense over node attributes and graph topology as the inner certified classifier and outer certified classifier, respectively. Now, a natural question is: is it possible to achieve certified defense in a different order, i.e., first achieve certified defense over the graph topology (as the inner classifier), and then realize certified defense over the node attributes (as the outer classifier)? Note that this is not the research focus of this paper, but we will provide insights about this question. In fact, it is also feasible to achieve certified defense in the reversed order compared with the approach presented in our paper. We provide an illustration in Figure˜5. We follow a similar setting to plot this figure as in Section 3.3. Specifically, in case (1), both $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$ lead to a positive outcome for $g$; in case (2), both $\bm{A}_{i,j}\oplus 0$ and $\bm{A}_{i,j}\oplus 1$ lead to a negative outcome. However, considering the Gaussian distribution around $\bm{X}_{i,j}$, samples will fall around case (1) with a much higher number compared with case (2). Hence, in this example, it would be reasonable to assume that the classifier with Bernoulli noise over graph topology (the inner certified classifier) will return 1 with a higher probability. This example illustrates how certification via a different order returns 1.
@@
However, such a formulation bears higher computational costs in calculating the certified budgets. The reason is that we are able to utilize a closed-form solution to calculate $\epsilon_{\bm{X}}$ based on a set of Gaussian noise and the corresponding output from the bias indicator function. However, based on a set of Bernoulli noise and the corresponding output from the bias indicator function, we will need to solve the optimization problem given in Theorem˜5 to calculate $\epsilon_{\bm{A}}$, which bears a higher time complexity than calculating $\epsilon_{\bm{X}}$. If we follow the strategy provided in Section 3.4 to calculate the inner and outer certification budgets, the certified budget of the inner certification will always be calculated multiple times, while the certified budget of the outer certification will only be calculated once. Considering the high computational cost of calculating $\epsilon_{\bm{A}}$, we thus argue that it is more efficient to realize the certification over graph topology as the outer certified classifier.
B.5. Time Complexity Analysis
@@
Experimental. We perform a study of running time, and we present the results in Table˜4. Specifically, we compare the running time of a successful certification under 30,000 random noise samples and a regular training-inference cycle with vanilla GCN. We observe that (1) although ELEGANT improves the computational cost compared with the vanilla GNN backbones, the running time remains acceptable; and (2) ELEGANT has less running time growth rate on larger datasets. For example, E-SAGE has around 10x running time on German Credit (a smaller dataset) while only around 4x on Credit Default (a larger dataset) compared to vanilla SAGE. Hence we argue that ELEGANT bears a high level of usability in terms of complexity and running time.
B.6. Additional Results on Different GNN Backbones & Baselines
@@
We perform additional experiments over two popular GNNs, including APPNP (KlicperaBG19) and GCNII (chen2020simple), to evaluate the generalization ability of ELEGANT onto different backbones. We present all numerical results in Table˜5 (in terms of accuracy), Table˜6 (in terms of fairness), and Table˜7 (in terms of FCR). We observe that ELEGANT achieves comparable utility, a superior level of fairness, and a large percentage of FCR. This verifies the satisfying usability of ELEGANT, which remains consistent with the paper.
@@
In addition, we provide a detailed fairness comparison between ELEGANT and robust GNNs from (jin2021node) and (wu2019adversarial) in Table˜8. We observe that the best performances still come from the GNNs equipped with ELEGANT on all datasets. Hence we argue that ELEGANT exhibits satisfying performance in usability, which remains consistent with the discussion in the paper.
@@
Why ELEGANT Improves Fairness? We note that improving fairness is a byproduct of ELEGANT, and our focus is to achieve certification over the fairness level of the prediction results. We now provide a detailed discussion about why fairness is improved here. First, existing works found that the distribution difference in the node attribute values and edge existence across different subgroups is a significant source of bias (dong2021edits; dai2021say; fan2021fair). However, adding noise on both node attributes and graph topology may reduce such distributional divergence and mitigate bias. Second, As mentioned in Section˜3.4, the proposed strategy to obtain the output predictions in ELEGANT is to select the fairest result among the output set $\hat{\mathcal{Y}}^{\prime}$, where each output is derived based on a sample $\bm{\Gamma}_{\bm{A}}^{\prime}\in\bar{\mathcal{A}}^{\prime}$ (i.e., $\mathrm{argmin}_{\hat{\bm{Y}}^{\prime}}\pi(\hat{\bm{Y}}^{\prime},\mathcal{V}_{\text{tst}})\;\;s.t.\;\hat{\bm{Y}}^{\prime}\in\hat{\mathcal{Y}}^{\prime}$). Such a strategy provides a large enough probability to achieve certification in light of Proposition 1. Meanwhile, we note that such a strategy also helps to significantly improve fairness since highly biased outputs are excluded.
B.7. Complementary Results
@@
We provide the results in terms of $\Delta_{EO}$ for Table˜2 in Table˜9, and we present the results of the baselines for Figure˜3 in Table˜10, Table˜11, Table˜12, and Table˜13. For Table˜9, we observe that ELEGANT does not constantly show a lower value of $\Delta_{EO}$. This is because the certification goal in Table˜2 is $\Delta_{SP}$ instead of $\Delta_{EO}$. In addition, we note that debiasing existing GNN models is not the goal of this paper. In addition, we provide the corresponding results in terms of accuracy for Figure˜4 in Table˜14 and Table˜15. We observe that although most performance remains stable, a stronger noise (i.e., larger $\sigma$ and smaller $\beta$) generally leads to worse but still comparable performance. This is consistent with the discussion in Section˜4.4, and this has been taken into consideration in the discussion of the parameter selection strategy in Section˜4.4.
@@
Table 4. Comparison of running time (in seconds) on different datasets using different methods.
@@
Table 5. Performance comparison of classification accuracy. Numbers are in percentage.
@@
Table 6. Comparison of fairness (measured with $\Delta_{SP}$). Numbers are in percentage.
@@
Table 7. Performance in FCR on different datasets and backbone GNNs. Numbers are in percentage.
@@
Table 8. Comparison of fairness (measured with $\Delta_{SP}$). Numbers are in percentage.
@@
(jin2021node) 14.8 $\pm$ 18.3 9.59 $\pm$ 0.65 3.84 $\pm$ 0.17
(wu2019adversarial) 3.66 $\pm$ 0.52 8.04 $\pm$ 2.97 7.10 $\pm$ 5.10
Table 9. The $\Delta_{EO}$ of Table˜2 in the paper. All numerical numbers are in percentage.
@@
Table 10. The results under ($2^{0}$, $10^{-1}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ Figure˜3. All numerical numbers are in percentage.
@@
Table 11. The results under ($2^{1}$, $10^{0}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ Figure˜3. All numerical numbers are in percentage.
@@
Table 12. The results under ($2^{2}$, $10^{1}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ for Figure˜3. All numerical numbers are in percentage.
@@
Table 13. The results under ($2^{3}$, $10^{2}$) in terms of node classification accuracy, AUC score, F1 score, $\Delta_{SP}$, and $\Delta_{EO}$ for Figure˜3. All numerical numbers are in percentage.
@@
Table 14. Classification accuracy in Figure˜4(a) with different settings. Numbers are in percentage.
@@
Table 15. Classification accuracy in Figure˜4(b) with different settings. Numbers are in percentage.
@@ Appendix C Additional Discussion
C.1. Why Certify A Classifier on top of An Optimized GNN?
@@
Table 16. Experimental results on Pokec-z and Pokec-n datasets.
@@
C.2. What Is the Difference Between the Attacking Performance of GNNs and the Fairness of GNNs?
C.3. Certification Without Considering the Binary Sensitive Attribute
C.4. How Do the Main Theoretical Findings Differ From Existing Works on Robustness Certification of GNNs on Regular Attacks?
Most existing works for robustness certification can only defend against attacks on either node attributes or graph structure. Due to the multi-modal input data of GNNs, existing works usually fail to handle the attacks over node attributes and graph structure at the same time. However, ELEGANT is able to defend against attacks over both data modalities. This necessitates using both continuous and discrete noises for smoothing and the analysis for joint certification in the span of the two input data modalities.
C.5. Difference with Existing Works
Here we mainly focus on discussing the difference between this work and (bojchevski2020efficient). We note that (1) the randomized smoothing technique adopted in (bojchevski2020efficient) is different from the proposed randomized smoothing approach on the graph topology in this paper and (2) the techniques in (bojchevski2020efficient) tackle a different problem from this paper. We elaborate on more details below. The techniques in (bojchevski2020efficient) are different from this paper. Although both randomized smoothing approaches are able to handle binary data, we note that the randomized smoothing approach proposed in (bojchevski2020efficient) is data-dependent. However, the proposed randomized smoothing approach in this paper is data-independent. We note that in practice, a data-independent approach enables practitioners to pre-generate noises, which enhances usability.
The studied problem in (bojchevski2020efficient) is different from this paper. Although the authors claimed to achieve a joint certificate for graph topology and node attributes in (bojchevski2020efficient), all node attributes are assumed to be binary, which can only be applied to cases where these attributes are constructed as bag-of-words representations (as mentioned in the second last paragraph in the Introduction of (bojchevski2020efficient)). However, in this work, we follow a more realistic setting where only graph topology is assumed to be binary while node attributes are considered as continuous. This makes the problem more difficult to handle, since different strategies should be adopted for different data modalities. In summary, compared with (bojchevski2020efficient), the problem studied in this paper is more realistic and more suitable for GNNs.
Based on the discussion above, we would like to note that no existing work can be directly adopted to tackle the studied problem in this paper, which further evidenced the novelty of this work.
C.6. Additional Experiments on Other Datasets
To further validate the performance of the proposed method, we also perform experiments with the same commonly used popular GNN backbone models (as Section 4.2) on two Pokec datasets, namely Pokec-z and Pokec-n. We present the experimental results in Table˜16, where all numerical numbers are in percentage. We observe that (1) the GNNs equipped with ELEGANT achieve comparable node classification accuracy; (2) the GNNs equipped with ELEGANT achieve consistently lower levels of bias; and (3) the values of the Fairness Certification Rate (FCR) for all GNNs equipped with ELEGANT exceed 90%, exhibiting satisfying usability. All three observations are consistent with the experimental results and the corresponding discussion presented in Section 4.2. Therefore, we argue that the effectiveness of ELEGANT is not determined by the dataset and is well generalizable over different datasets.
C.7. Scalability of ELEGANT